OpenStudy (anonymous):

how does 5sin111.8/10=27.7?

OpenStudy (anonymous):

\[5\sin111.8\over10\]

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

What makes you think it does?

OpenStudy (anonymous):

its the answer in the book

OpenStudy (whpalmer4):

Perhaps the left hand side of that equation is not correct :-)

OpenStudy (anonymous):

nope. im studying for an exam and it gives you the solution to some problems to show you how to do it. and im going through the steps and that what it says

OpenStudy (whpalmer4):

Show me the entire problem, please.

OpenStudy (anonymous):

here is a photo if you dont beleive me ill write out the entire problem one sec

OpenStudy (anonymous):



OpenStudy (whpalmer4):

That's a completely different statement than you gave us, even accounting for the picture being reversed :-) \[\sin A = \frac{5 \sin 111.8^\circ}{10}\rightarrow \text{angle }A = 27.7^\circ\] \[\sin A = \frac{5*0.928486}{10} \approx 0.464243\]\[A = \sin^{-1} 0.464243 \approx 0.48278 \text{ radians} = 27.6612^\circ\]

OpenStudy (anonymous):

ooooh okay so inverse sine.

OpenStudy (anonymous):

thanks :)

OpenStudy (whpalmer4):

you're welcome!